# Spontaneous time-reversal symmetry breaking by disorder in superconductors

### v1

@@ 3.3 Enhanced RKKY exchange coupling and slowing-down of magnetic fluctuations by superconductivity
The first scenario (A) was largely discussed in Sec. 3.2, there exemplified through models relevant for cuprates, i.e. -wave superconductivity in the presence of nonmagnetic impurities and electronic correlations. This mechanism is likely much more broadly applicable since most unconventional superconductors allow for in-gap bound states from nonmagnetic disorder [111, 82, 112]. This causes enhanced local spin susceptibilities and associated tendency to slow down magnetic fluctuations locally or even pin magnetic moments [80]. This mechanism may well also be at play near e.g. sample edges or various forms of dislocations [52, 53]. A related property may happen near sample edges when zero-energy Andreev bound states generated by sign-changes of the gap function split by spontaneous generation of edge supercurrents [74, 76, 78].
@@
Finally we discuss scenario (C) where a distribution of disordered static magnetic moments are already present in the normal state. Thus in this case, there is strictly speaking no spontaneous TRSB in the superconducting state only. We do not discuss in detail the origin of such moments, but they could simply arise from weakly coupled magnetic impurities, or small spin-polarized regions that have crossed into a nearby magnetic phase due to chemical disorder. It might also be due to the mechanism discussed in Sec. 3.2 if the "normal state" exhibits pseudo-gap behavior. We assume that the magnetism in the normal state is too weak to be detected by standard probes, for example due to a tiny volume fraction and/or short-range nature. A TRSB signal may be detected, however, in cases where the volume fraction grows and/or the moments couple more strongly, and the magnetic quasi-order becomes longer ranged as a result upon entering the superconducting phase. The latter may arise from the mechanism discussed under (B). It might also take place via new "stepping stone" magnetic islands produced by nonmagnetic disorder in the superconducting state via the mechanism discussed in Sec. 3.2. An enhanced volume fraction can take place in cases where another competing order is in the game. If local disorder-induced magnetism is limited by orbital order or a CDW phase, for example, then superconductivity may suppress these competing orders and thereby alleviate the magnetism. Clearly, these are merely possibilities for observing TRSB below , and need further detailed investigations for specific materials.
@@ 5 Summary and Outlook
Naturally an important question relates to the relevance of these mechanisms to the observation of TRSB below in currently discussed unconventional superconducting materials. In the above, we provided a detailed discussion of the cuprates and the possibility of superconductivity-enhanced spin-polarization. Two other materials of significant current interest are UTe and SrRuO. For these compounds, two-component condensates are widely discussed, but inconsistent with specific heat measurements, also under uniaxial strain featuring no signs of double transitions [32, 33, 34, 35]. This points to single-component superconductivity which has received recent support from a number of other probes as well [144, 136, 108, 145], but leaves open the question of what causes a TRSB signal in some measurements. We have surveyed several possible answers to this question from the perspective of induced fields within the superconducting state from spatial inhomogeneity.

### v2

@@ 3.3 Enhanced RKKY exchange coupling and slowing-down of magnetic fluctuations by superconductivity
The first scenario (A) was largely discussed in Sec. 3.2, there exemplified through models relevant for cuprates, i.e. -wave superconductivity in the presence of nonmagnetic impurities and electronic correlations. This mechanism is likely much more broadly applicable since most unconventional superconductors allow for in-gap bound states from nonmagnetic disorder [123, 90, 124]. This causes enhanced local spin susceptibilities and associated tendency to slow down magnetic fluctuations locally or even pin magnetic moments [88]. This mechanism may well also be at play near e.g. sample edges or various forms of dislocations [57, 58, 125, 126]. A related property may happen near sample edges when zero-energy Andreev bound states generated by sign-changes of the gap function split by spontaneous generation of edge supercurrents [82, 84, 86].
@@
Finally we discuss scenario (C) where a distribution of disordered static magnetic moments are already present in the normal state. Thus in this case, strictly speaking, TRSB is not restricted to the superconducting phase below . We do not discuss in detail the origin of such moments, but they could simply arise from weakly coupled magnetic impurities, or small spin-polarized regions that have crossed into a nearby magnetic phase due to chemical disorder. It might also be due to the mechanism discussed in Sec. 3.2 if the "normal state" exhibits pseudo-gap behavior. We assume that the magnetism in the normal state is too weak to be detected by standard probes, for example due to a tiny volume fraction and/or short-range nature. A TRSB signal may be detected, however, in cases where the volume fraction grows and/or the moments couple more strongly, and the magnetic quasi-order becomes longer ranged as a result upon entering the superconducting phase [139, 116]. The latter may arise from the mechanism discussed under (B). It might also take place via new "stepping stone" magnetic islands produced by nonmagnetic disorder in the superconducting state via the mechanism discussed in Sec. 3.2. An enhanced volume fraction can take place in cases where another competing order is in the game. If local disorder-induced magnetism is limited by orbital order or a CDW phase, for example, then superconductivity may suppress these competing orders and thereby alleviate the magnetism. Clearly, these are merely possibilities for observing TRSB below , and need further detailed investigations for specific materials.
@@ 5 Summary and Outlook
Naturally an important question relates to the relevance of these mechanisms to the observation of TRSB below in currently discussed unconventional superconducting materials. In the above, we provided a detailed discussion of the cuprates and the possibility of superconductivity-enhanced spin-polarization. Two other materials of significant current interest are UTe2 and Sr2RuO4. For these compounds, two-component condensates are widely discussed, but inconsistent with specific heat measurements, also under uniaxial strain featuring no signs of double transitions [36, 37, 38, 39]. This points to single-component superconductivity which has received recent support from a number of other probes as well [158, 150, 117, 159], but leaves open the question of what causes a TRSB signal in some measurements. We have surveyed several possible answers to this question from the perspective of induced fields within the superconducting state from spatial inhomogeneity.